Sulfuric acid (H2SO4), a highly corrosive strong mineral acid, reacts with Cu according to this equation: 2H2SO4 + Cu → CuSO4 + 2H2O + SO2 How many grams of sulfur dioxide (SO2) are created by this reaction if 4.78 mol of Cu react with the H2SO4?


1
Expert's answer
2021-11-11T11:03:01-0500

CU:SO2= 1:1

= 4.78

H2SO4:SO2= 2:1

= 4.78/2= 2.39

4.78+2.39= 7.17moles